Shift Summary for DAY shift, 9 Aug 99

Julie Whitmore (jaws@fnpxsr)
Wed, 11 Aug 1999 16:43:24 CDT

Shift Summary for DAY shift, 9 Aug 99

On shift: Val Prasad, Brent VanDevender
Also present: Julie, Sydney

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 3.5
Hours of E832 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: beam down to fix rf problems,tape change

==================================================================

Runs ended this shift:
14151 TAPES QK*240-242,QKB084
14152 New Tapes : QK*243-245,QKB085
14153 (laser scan)

========================================================================
Accesses during shift:
none

========================================================================
Overview of shift:
Beam at low intensity due to rf problems until almost 1100. After about an
hour of downtime beam returns at normal intensity and remains throughout shift.
CsI problems (drifting gain and channel out of illumination).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14151 7:30 Take over from owl shift
10:08 run stopped (tapes full) : QK*240-242,QKB084

14152 10:15 began new run : Tapes QK*243-245,QKB085
10:48 beam disappears (TeV rf problems)
11:14 run stopped

14153 11:15 began laser scan
11:52 beam returns; laser scan aborted after 2.5 passes

14154 11:53 new run started
11:57 TeV rf repairs are complete and beam is back at
~6E12
12:41 spill 48: CsI channel 1137 drops out of illumination
14:30 channel 116 has large (+/- 10 channels) pedestal
oscillations. Unsparsified it following
Elizabeth's suggestion (effective next run).
If Run Control prompts FERA Pedestal update
just ignore for now.

===================================================================

Follow up notes on previous shift:

===================================================================
Items to follow up on next shift:
MU3Y ADC may require High Voltage Adjustment on next access.
Cosmic pedestal T17 needs to be checked.
DO NOT UPDATE FERA Pedestals when you start the next run.